OpenStudy (anonymous):

PLZ HELP!!!!!! Will give medals. Combine and simplify the following radical expression. ((2)/(3^3√5)) Combine and simplify the following radical expression. ((√5) + (√3)/(√3) - (√5))

OpenStudy (anonymous):

Also (2*√x+2) - (√x-3) = (√x+9)

OpenStudy (isaiah.feynman):

\[\frac{ 2 }{3 ^{3\sqrt{5}} }\]There is nothing else to do to that.

OpenStudy (anonymous):

i think he is right

OpenStudy (anonymous):

it has to be a fraction with a radical on top.

OpenStudy (anonymous):

ok let me see

OpenStudy (anonymous):

(233√5)

OpenStudy (anonymous):

Over what?

OpenStudy (anonymous):

wait a sec

OpenStudy (anonymous):

((√5)+√3√3−(√5))

OpenStudy (anonymous):

might be wrong

OpenStudy (anonymous):

but pretty sure it aint